# Unlocking the Capabilities of Vision-Language Models for Generalizable and Explainable Deepfake Detection

### v1

@@ Abstract
This paper presents a novel deepfake detection framework based on Large Vision-Language Model (LVLM) that integrates external knowledge to improve generalization and explainability. Previous deepfake detection methods often rely on data or feature augmentation to increase diversity but lack the incorporation of human knowledge, which limits their ability to generalize across different forgery types and unseen manipulations. To address these limitations, we first train a generalized detector by introducing external knowledge, then fine-tune the LVLM to enable multi-turn dialogue capabilities. Our framework includes a Knowledge-guided Forgery Detector, an image encoder, and a Large Language Model (LLM). The image encoder extracts visual prompt embeddings from images, while the LLM receives visual and question prompt embeddings for inference. The Knowledge-guided Forgery Detection Module is used to calculate correlations between image features and pristine/deepfake class embeddings, enabling forgery classification and localization. The outputs from these components are processed by a Forgery Prompt Learner to construct forgery prompt embeddings. Finally, we feed these prompt embeddings into the LLM to generate textual detection response to assist judgment. Extensive experiments on multiple benchmarks, including FF++, CDF2, DFD, DFDCP, and DFDC, demonstrate that our scheme surpasses state-of-the-art methods in generalization performance, while also supporting multi-turn dialogue capabilities.
Machine Learning, ICML
@@ 1 Introduction
Current deepfake detection research primarily focuses on enhancing generalization and accuracy for identifying previously unseen forgeries. Various approaches have been explored, including feature consistency analysis (Dong et al., 2022; Zhao et al., 2021b; Liu et al., 2023), reconstruction error techniques (Cao et al., 2022; Chen et al., 2021; Shi et al., 2023), and methods focusing on visual artifacts (Wang et al., 2023b; Sun et al., 2023; Nguyen et al., 2024). While these methods effectively leverage image-level or feature-level anomalies, they often struggle to fulfill interpretability requirements in real-world scenarios. In judicial settings, however, existing deepfake detection systems must also support conversational capabilities to address inquiries from judges and skeptics. Consequently, the key challenge lies in providing sufficient interpretability without compromising generalization performance.
Thus, our objective is to develop a generalized and explainable deepfake detection scheme. Large Vision-Language Models (LVLMs) presents a promising leverage to this issue. Pre-trained on extensive and diverse datasets, LVLMs capture vast amounts of knowledge about natural objects, providing significant potential to improve generalization in identifying manipulated content. Furthermore, LVLMs exhibit capabilities for effective learning in low-sample scenarios, making them particularly suitable for tasks with limited annotated data (Xu et al., 2024; Huang et al., 2024). A LVLM typically uses an image encoder to extract image features, which are then combined with text prompts and input into the Large Language Model (LLM) to generate responses. For instance, inputting an image along with a prompt like “This is a facial image designed for deepfake detection, and it should not exhibit any localized color discrepancies. Is this a deepfake image?” allows the LVLM to assess potential manipulations. However, existing LVLMs are primarily optimized for general image understanding tasks and may not effectively capture the fine-grained features required for deepfake detection. Directly performing fine-tuning presents challenges, as the LVLM may struggle to interpret specialized terms like “visual artifacts” or “inconsistencies” as intended in the context of forgery detection. Therefore, it is crucial to design fine-grained prompt embeddings to facilitate LVLM training.
To address these challenges, we propose a novel LVLM-based deepfake detection framework under the guidance of pretrained knowledge of LVLM to make generalizable and explainable detection (Lester et al., 2021; Liu et al., 2022). As shown in Figure 1, we first integrate external knowledge to train a generalizable detector, and then incorporate its features into LLM to generate responses. Specifically, the framework comprises two key stages: Knowledge-guided Forgery Detector Training and LLM Prompt Tuning. In the Knowledge-Guided Forgery Detector Training phase, we aim to train a high-precision deepfake detector. Leveraging a pre-trained multimodal model, we extract both image content features and textual features from descriptions of real and fake images. By calculating the correlation between these image features and learnable real/fake class embeddings, we generate consistency maps that represent the alignment between visual content and textual descriptions. These maps are subsequently processed by the Forgery Locator and Classifier to produce a forgery segmentation map and forgery score. Subsequently, in the LLM Prompt Tuning phase, we incorporate deepfake detection knowledge into the LLM to generate forgery detection descriptions. To ensure accurate deepfake detection, we train the LVLM using simulated forgery image-text pairs specifically tailored for this task. Our contributions are summarized as follows:
@@
  • We introduce a Knowledge-guided Forgery Detector to generate forgery consistency maps to align image features with textual descriptions of both pristine and deepfake images for enhanced generalization.
  • Extensive experiments on multiple benchmarks, including FF++, CDF1, CDF2, DFD, DFDCP, and DFDC, demonstrate that our scheme outperforms existing methods in generalization performance, with the added capability of supporting multi-turn dialogue.
@@ 2 Related Works
Deepfake Detection Methods: Conventional classification architectures have achieved significant success in detecting forgery clues in early deepfake methods. Various strategies, such as data augmentation (Li et al., 2020a; Shiohara & Yamasaki, 2022; Nguyen et al., 2024), feature consistency analysis (Zhao et al., 2021b; Yan et al., 2023), and frequency domain anomaly detection (Jeong et al., 2022; Liu et al., 2021; Wang et al., 2023a), have been explored in recent years to enhance the generalization of deepfake detection models. While these methods achieve a high detection accuracy, they primarily rely on data augmentation or enhanced feature extraction (Yan et al., 2024), and often neglect the integration of external human knowledge. However, many deepfake characteristics are embedded in human knowledge, which is challenging to capture through data or feature augmentation alone. This limitation significantly constrains the generalization capabilities of existing algorithms. In this paper, we propose an LVLM-based deepfake detection framework that aligns image features with real/fake descriptions to enhance the model’s capacity to detect unseen deepfakes.
@@ 3 Proposed Method
Our objective is to enable LVLMs to accurately distinguish between real and fake faces. Although LVLMs are trained on large-scale datasets, they are primarily trained for general image understanding tasks and often lack the sensitivity for detecting forgery details. To address this limitation, we propose a novel LVLM-based deepfake detection framework that enhances sensitivity to deepfake artifacts through constructing fine-grained forgery prompts. As shown in Figure 2, our scheme builds upon a conventional LVLM framework, which comprises an image encoder, a projector, and an LLM. The image encoder extracts content features from input images, which are subsequently transformed into visual prompt embeddings $E_{visual}$ by the projector. Additionally, user queries are tokenized into question prompt embeddings $E_{question}$. To train the model for forgery detection, we employ a two-stage process. In the first stage, we train a Knowledge-guided Forgery Detector (KFD) to perform forgery detection and localization by calculating the correlation between image content features and deepfake descriptions. This stage ensures that the KFD can effectively classify and localize forgery artifacts by learning fine-grained visual-text correlations. In the second stage, we perform LLM Prompt Tuning to integrate the KFD knowledge into the LVLM framework. Specifically, we design a Forgery Prompt Learner to convert forgery-related features into forgery prompt embeddings. These embeddings, along with the question and visual prompt embeddings, are then fed into the LLM to generate a textual detection result. By employing prompt tuning and alternating training with pre-trained data, our scheme ensures accurate deepfake detection while preserving the LLM’s multi-turn dialogue capability.
Figure 3: Overview of the Knowledge-guided Forgery Detector. It uses a pretrained image encoder and text encoder to extract features from both modalities, and computes their correlation to produce consistency maps. These maps are further processed by a forgery classifier and a locator, yielding a forgery score and a forgery segmentation map. The Reference based Optimization process is applied exclusively during the training phase to enhance the robustness of extracted features.
@@ 3.1 Knowledge-guided Forgery Detector
Forgery Visual-Text Alignment: To acquire forgery detection-related knowledge, inspired by (Jeong et al., 2023), we align image content features with predefined text description features to obtain fine-grained forgery features. This process is illustrated in Figure 3. Specifically, this process involves a pretrained image encoder, a pretrained text encoder, and a linear layer. Both the image and text encoders are sourced from ImageBind, a large-scale multimodal pre-trained model with extensive cross-modal knowledge. We first define real and fake image descriptions, $D_{real}$ and $D_{fake}$, and augment them with learnable context to enable the Text Encoder to extract task-relevant textual features, $F_{text}\in\mathbb{R}^{2\times C_{text}}$. For the visual features, we select $l$ layers from the image encoder and obtain the intermediate features extracted by each selected layer. The extracted features are then processed by a linear layer to generate visual features $F_{vis}^{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{text}}$, where $i$ indicates the $i$-th layer. The similarity maps between visual features and textual features are calculated and concatenated as consistency maps. The formula for computing the consistency maps is as follows:
$C_{text}=\{F_{vis}^{i}F_{text}^{T}\}.$ (1)
@@
Additionally, to optimize the extracted image features, we compute the similarity between the features of a reference image (pristine image) $F_{vref}^{i}$ and the input image features $F_{vis}^{i}$. This similarity optimization enhances the robustness of the features extracted by the image encoder. Note that reference images are used only during training and are not employed at test time. The similarity is calculated as:
$C_{ref}=\{Cos(F_{vref}^{i},F_{vis}^{i})\}.$ (2)
@@
Forgery Locator and Classifier: To enhance the model’s sensitivity to deepfake content, we introduce a Forgery Locator and a Forgery Classifier to locate forgery areas and classify pristine and deepfake images. As shown in Figure 3, the Forgery Locator consists of three branches. Each branch applies down-sampling and up-sampling operations on the corresponding consistency maps, followed by concatenation and convolution operations to generate the segmentation map. The Forgery Classifier also consists of three branches. The three feature maps after convolution are down-sampledbefore they are concatenated. After that, we use two fully connected layers to calculate the probability of the image being real or fake. Here, we use Dice Loss to improve the accuracy of forgery segmentation. We further enhance the robustness of extracted forgery features by calculating the matching degree between segmentation maps for the textual consistency map ($C_{text}$) and the reference map ($C_{ref}$). Both maps are expected to accurately localize the forged regions. The localization loss is formulated as follows:
$\mathcal{L}_{loc}=Dice(\phi(C_{text}),gt)+\lambda Dice(\phi(C_{ref}),gt),$ (3)
@@
where $\phi$ is the Forgery Locator, and $gt$ is the ground truth mask. Dice loss measures the overlap between the predicted segmentation and the ground truth. $\lambda$ is the loss weight for balancing these two losses.
@@ 3.2 Forgery Prompt Learner and LLM
Figure 4: Architecture of the Forgery Prompt Learner.
Forgery Prompt Learner: To effectively convert forgery-related features into inputs for the LLM, we propose a Forgery Prompt Learner to transform the forgery localization map, the forgery score, and the consistency maps into forgery prompt embeddings. At the same time, we add a learnable prompt embeddings for the forgery prompt learner to incorporate extra information for the deepfake detection task. As shown in Figure 4, the Forgery Prompt Learner consists of two convolutional neural networks, one fully connected layer, and the learnable prompt embeddings $E_{\emph{base}}\in\mathbb{R}^{n_{1}\times C_{\emph{emb}}}$. Specifically, the two convolutional networks transform the forgery localization map and consistency maps into vector representations, $E_{\emph{loc}}\in\mathbb{R}^{n_{2}\times C_{\emph{emb}}}$ and $E_{\emph{cons}}\in\mathbb{R}^{n_{3}\times C_{\emph{emb}}}$, respectively. The forgery score is expanded into $E_{\emph{cls}}\in\mathbb{R}^{1\times C_{\emph{emb}}}$. These embeddings are concatenated and fed into the convolution layer to generate the forgery prompt embeddings $E_{\emph{forgery}}\in\mathbb{R}^{n_{f}\times C_{\emph{emb}}}$. Finally, the forgery prompt embeddings, visual prompt embeddings, and question prompt embeddings are input into the LLM.
LLM: The LLM processes the prompt embeddings to interpret the context and accurately identify forged regions. By integrating visual details (from $E_{\emph{forgery}}$ and $E_{\emph{visual}}$) with user queries (from $E_{\emph{question}}$), the LLM produces responses that provide forgery detection judgments and precisely localize manipulated regions (e.g., eyes, mouth). To optimize model performance, we employ prompt tuning on the LLM using simulated image-text pairs specifically designed for deepfake detection. To ensure the LLM generates accurate responses, we use cross-entropy loss to quantify the discrepancy between the predicted response and the target label. The formula is defined as follows:
@@
where $\hat{y_{j}}$ and $y_{j}$ are the predicted probability , and the corresponding ground truth label, respectively for the $j$-th token in the sequence.
@@ 3.3 Data for LLM Prompt Tuning
Forgery Data Simulation: We intend for the LLM to identify pristine and deepfake images, while also locating the forged regions. This requires training on image-text pairs specifically depicting manipulated areas, which are currently unavailable. To address this gap, we draw on techniques from SBI (Shiohara & Yamasaki, 2022) to construct image-text pairs using existing real images. First, we generate facial landmarks from a real image $I_{\emph{real}}$, then randomly select 1 to 3 regions (e.g., the nose, mouth, or eyes) as target forgery areas. We apply a slight affine transformation to the real image, resulting in an affine-transformed image $I_{\emph{affine}}$. The original real image is used as the background (target face), and the affine-transformed image serves as the foreground (source face). Following the approach in (Nguyen et al., 2024), we apply Poisson blending to combine the foreground and background images. The blending process is defined as follows:
@@
Methods Intra-dataset Cross-dataset FF++ CDF2 DFD DFDC CDF1 DFDCP AUC AP AUC AP AUC AP AUC AP AUC AP AUC AP Xception (Rossler et al., 2019) CVPR’19 97.23 96.44 81.65 88.74 89.75 - - - 80.98 90.00 69.90 81.95 FaceXRay+BI (Li et al., 2020a) CVPR’20 - - 79.50 - 95.40 93.34 - - 80.58 73.33 80.92 72.65 F3Net (Qian et al., 2020) ECCV’20 98.20 96.17 78.88 86.23 93.33 97.70 71.77 71.99 81.11 89.61 73.50 79.30 Multiattentional (Zhao et al., 2021a) CVPR’21 - - 68.26 - 92.95 - - - - - 63.02 - SPSL (Liu et al., 2021) CVPR’21 96.91 89.37 79.86 87.83 96.12 98.20 66.16 71.13 85.02 92.17 75.86 82.64 PCL+I2G (Zhao et al., 2021b) CVPR’21 99.11 - 90.03 - 99.07 - - - - - 74.27 - RECCE (Cao et al., 2022) CVPR’22 99.32 97.25 82.31 88.26 98.26 98.40 69.58 69.98 81.49 89.04 71.49 77.04 SBI (Shiohara & Yamasaki, 2022) CVPR’22 99.15 99.15 93.82 92.99 96.32 96.13 74.47 74.09 93.44 93.77 90.95 85.98 SFDG (Wang et al., 2023a) CVPR’23 - - 75.83 - 88.00 - - - - - 73.63 - AltFreezing (Wang et al., 2023b) CVPR’23 93.81 98.74 89.50 87.41 98.50 94.59 64.75 67.52 88.48 92.79 64.05 76.22 CADDM (Dong et al., 2023) CVPR’23 99.70 - 93.88 - 99.03 - 73.85 - 85.68 - 74.19 - UCF (Yan et al., 2023) CVPR’23 98.69 97.99 83.73 90.10 94.50 98.04 75.11 74.76 86.08 91.78 80.50 77.16 TALL (Xu et al., 2023) ICCV’23 99.87 - 90.79 - - - 76.78 - - - - - LSDA (Yan et al., 2024) CVPR’24 - - 91.10 - - - 77.00 - - - - - LAANet (Nguyen et al., 2024) CVPR’24 99.96 - 95.40 97.64 99.51 99.40 - - - - 86.94 97.70 Ours 99.53 99.54 94.71 93.59 99.64 99.68 79.12 77.69 97.62 97.67 91.81 88.26
@@
Table 1: Intra-dataset and cross-dataset evaluation of KFD compared with existing deepfake detection methods. The highest AUC and AP scores among all methods are highlighted in bold, while the second-highest scores are underlined. All experimental results are sourced from the original papers or publicly available code repositories.
@@
Question and Answer Content: Training an LVLM requires a large number of visual question-answer pairs. Therefore, we construct corresponding text queries for each image. To ensure compatibility with the deepfake detection task, we first include a background description in each query, for example: “This is a facial image designed for deepfake detection, and it should not exhibit any localized color discrepancies or evident signs of splicing.”, which can be regarded as a kind of human prior knowledge. Additionally, we incorporate the prediction results from the Knowledge-Guided Forgery Detector (KFD) into the prompt, such as According to KFD prediction, the forgery score is 0.93. We then ask a question related to the content of the image, such as “Is this a deepfake image?” The LVLM’s response states whether a forgery is present in the image and where the forgery area is. For instance, “Yes. This is a deepfake image, and the artifact is at the center face of the image.” Here, the forgery regions are defined according to those selected during forgery data simulation. By defining both queries and responses, we can train the LVLM to distinguish between pristine and deepfake images. The prompt format input to the LVLM follows this format:
@@
where $E_{visual}$ represents visual prompt embeddings, $E_{{forgery}}$ denotes the forgery prompt embeddings learned by the Forgery Prompt Learner, KFD result indicates the forgery score, and Task description provides a textual description of deepfake detection task.

### v2

# Unlocking the Capabilities of Large Vision-Language Models for Generalizable and Explainable Deepfake Detection

@@ Abstract
Current Large Vision-Language Models (LVLMs) have demonstrated remarkable capabilities in understanding multimodal data, but their potential remains underexplored for deepfake detection due to the misalignment of their knowledge and forensics patterns. To this end, we present a novel framework that unlocks LVLMs’ potential capabilities for deepfake detection. Our framework includes a Knowledge-guided Forgery Detector (KFD), a Forgery Prompt Learner (FPL), and a Large Language Model (LLM). The KFD is used to calculate correlations between image features and pristine/deepfake image description embeddings, enabling forgery classification and localization. The outputs of the KFD are subsequently processed by the Forgery Prompt Learner to construct fine-grained forgery prompt embeddings. These embeddings, along with visual and question prompt embeddings, are fed into the LLM to generate textual detection responses. Extensive experiments on multiple benchmarks, including FF++, CDF2, DFD, DFDCP, DFDC, and DF40, demonstrate that our scheme surpasses state-of-the-art methods in generalization performance, while also supporting multi-turn dialogue capabilities.
large vision language model, deepfake detection, generalization
@@ 1 Introduction
Large Vision-Language Models (LVLMs) present a promising leverage to this issue. Pre-trained on extensive and diverse datasets, LVLMs capture vast amounts of knowledge about natural objects, providing significant potential to improve generalization in identifying manipulated content. An LVLM typically uses an image encoder to extract image features, which are then combined with text prompts and input into the Large Language Model (LLM) to generate responses. For instance, inputting an image along with a prompt like “This is a facial image designed for deepfake detection, and it should not exhibit any localized color discrepancies or evident signs of splicing. Is this a deepfake image?” allows the LVLM to assess potential manipulations. However, existing LVLMs are primarily optimized for general image understanding tasks and may not effectively capture the fine-grained features required for deepfake detection. Directly performing fine-tuning presents challenges, as the LVLM may struggle to interpret specialized terms like “color discrepancies” or “visual artifacts” as intended in the context of forgery detection. Therefore, it is crucial to design fine-grained prompt embeddings to facilitate LVLM training.
@@
In this paper, we aim to unlock the capabilities of Large Vision-Language Models for deepfake detection tasks. Humans naturally use specific descriptors, such as subtle visual artifacts, localized lighting inconsistencies, and overly smoothed textures, to characterize manipulated content. However, these features are difficult to accurately replicate by data simulation or feature augmentation alone, limiting the ability of current methods to fully interpret manipulated content (Zhang et al., 2024). To address this limitation, we propose to explore the correction between images and textual descriptions to assist deepfake detection. As shown in Figure 1, we propose a novel LVLM-based deepfake detection framework under the guidance of pretrained knowledge to perform generalizable and explainable detection (Lester et al., 2021; Liu et al., 2022). We first integrate external knowledge to train a forgery detector, and then incorporate its features into LLM to generate responses. Specifically, the framework comprises two key stages: Knowledge-guided Forgery Detector Training and LLM Prompt Tuning. In the Knowledge-guided Forgery Detector Training phase, we aim to train a high-precision deepfake detector. Leveraging pre-trained multimodal encoders, we extract image features from images and textual features from descriptions of pristine and forged images. By calculating the correlation between these image features and description text embeddings, we generate consistency maps that represent the alignment between visual content and textual descriptions. These maps are subsequently processed by the Forgery Locator and Classifier to produce a forgery segmentation map and forgery score. Subsequently, in the LLM Prompt Tuning phase, we incorporate deepfake detection knowledge into the LLM to generate forgery detection descriptions. To ensure accurate deepfake detection, we train the LVLM using simulated forgery image-text pairs specifically tailored for this task. Our contributions are summarized as follows:
@@
  • We introduce a Knowledge-guided Forgery Detector to generate forgery consistency maps to align image features with description text embeddings of both pristine and deepfake images for enhanced generalization.
  • Extensive experiments on multiple benchmarks, including FF++, CDF1, CDF2, DFD, DFDCP, DFDC, and DF40, demonstrate that our scheme outperforms existing methods in generalization performance, with the added capability of supporting multi-turn dialogue.
@@ 2 Related Works
Deepfake Detection Methods: Conventional classification architectures have achieved significant success in detecting forgery clues in early deepfake content. Various strategies, such as data augmentation (Li et al., 2020a; Shiohara & Yamasaki, 2022; Nguyen et al., 2024), feature consistency analysis (Zhao et al., 2021b; Yan et al., 2023), and frequency domain anomaly detection (Jeong et al., 2022; Liu et al., 2021; Wang et al., 2023a), have been explored in recent years to enhance the generalization of deepfake detection models. While these methods achieve a high detection accuracy, they primarily rely on data augmentation or enhanced feature extraction (Yan et al., 2024), and often neglect the integration of external human knowledge. However, many deepfake characteristics are embedded in human knowledge, which is challenging to capture through data or feature augmentation alone. This limitation significantly constrains the generalization capabilities of existing algorithms. In this paper, we propose an LVLM-based deepfake detection framework that aligns image features with real/fake descriptions to enhance the model’s capacity to detect unseen deepfakes.
@@ 3 Proposed Method
Our objective is to enable LVLMs to accurately distinguish between real and fake faces. Although LVLMs are trained on large-scale datasets, they are primarily trained for general image understanding tasks and often lack the sensitivity for detecting forgery details. To address this limitation, we propose a novel LVLM-based deepfake detection framework that enhances sensitivity to deepfake artifacts through constructing fine-grained forgery prompts. As shown in Figure 2, our scheme builds upon a conventional LVLM framework, which comprises an image encoder, a projector, and an LLM. The image encoder extracts content features from input images, which are subsequently transformed into visual prompt embeddings $E_{visual}$ by the projector. Additionally, user queries are encoded as question prompt embeddings $E_{question}$. To train the model for forgery detection, we employ a two-stage process. In the first stage, we train a Knowledge-guided Forgery Detector (KFD) to perform forgery detection and localization by calculating the correlation between image content features and descriptions of pristine/deepfake images. This stage ensures that the KFD can effectively classify and localize forgery artifacts by learning fine-grained visual-text correlations. In the second stage, we perform LLM Prompt Tuning to integrate the KFD knowledge into the LVLM framework. Specifically, we design a Forgery Prompt Learner to convert forgery-related features into forgery prompt embeddings. These embeddings, along with the question and visual prompt embeddings, are then fed into the LLM to generate a textual detection result. To further enhance the interpretability of our model, we adopt an alternating training strategy using both deepfake datasets and general Visual Question Answering (VQA) datasets. This enables the model to accurately detect deepfakes while retaining multi-turn dialogue capabilities.
@@ 3.1 Knowledge-guided Forgery Detector
Forgery Visual-text Alignment: To acquire forgery detection-related knowledge, inspired by (Jeong et al., 2023), we align image content features with predefined text description features to obtain fine-grained forgery features. This process is illustrated in Figure 3. Specifically, this process involves a pretrained image encoder and a pretrained text encoder. Both the image and text encoders are sourced from ImageBind (Girdhar et al., 2023), a large-scale multimodal pre-trained model with extensive cross-modal knowledge.
Figure 3: Overview of the Knowledge-guided Forgery Detector. It computes correlation between image and text modalities to assist forgery classification and localization. The Reference based Optimization process is applied exclusively during the training phase to enhance the robustness of extracted features.
We first define real and fake image descriptions ($D_{real}$ and $D_{fake}$) and utilize a text encoder to extract their semantic features. These features are concatenated with a learnable embedding to obtain the task-specific text embedding $F_{text}\in\mathbb{R}^{2\times C_{text}}$, where $C_{text}$ denotes the dimensionality of the text embedding. For the visual features, we select $l$ layers from the image encoder and obtain the intermediate features extracted by each selected layer. The extracted features are then processed by linear layers to generate visual features $F_{vis}^{i}\in\mathbb{R}^{H_{i}\times W_{i}\times C_{text}}$, where $i$ indicates the $i$-th layer. The similarity maps between visual features and textual features are calculated and concatenated as consistency maps. The formula for computing the consistency maps is as follows:
$\rho_{text}=\{F_{vis}^{i}F_{text}^{\top}\}.$ (1)
@@
Additionally, to optimize the extracted image features, we compute the cosine similarity between the features of a reference image (pristine image) $F^{i}_{vref}$ and the input image features $F^{i}_{vis}$. This similarity optimization enhances the robustness of the features extracted by the image encoder. Note that reference images are used only during training and are not employed at inference time. The similarity is calculated as:
$\rho_{ref}=\{Cos(F_{vref}^{i},F_{vis}^{i})\}.$ (2)
@@
Forgery Locator and Classifier: To enhance the model’s sensitivity to deepfake content, we introduce a Forgery Locator and a Forgery Classifier to locate forgery areas and classify pristine and deepfake images. The Forgery Locator consists of three branches. Each branch performs down-sampling and up-sampling operations on the corresponding consistency maps, followed by interpolation, concatenation, and convolutional layers to generate the segmentation map. The Forgery Classifier also consists of three branches. The three feature maps are first processed by convolution and pooling operations, and then concatenated to form a unified feature representation. After that, we use two fully connected layers to calculate the probability of the image being real or fake. Here, we use Dice Loss to improve the accuracy of forgery segmentation. We further enhance the robustness of extracted forgery features by optimizing the matching degree between the textual consistency map ($\rho_{text}$) and the reference consistency map ($\rho_{ref}$). Both maps are expected to accurately localize the forged regions. The localization loss is formulated as follows:
$\mathcal{L}_{loc}=Dice(\phi(\rho_{text}),gt)+\lambda Dice(\phi(\rho_{ref}),gt),$ (3)
@@
where $\phi$ is the Forgery Locator, and $gt$ is the ground truth mask. Dice loss optimizes the overlap between the predicted segmentation and the ground truth. $\lambda$ is the loss weight for balancing these two losses.
@@ 3.2 Forgery Prompt Learner and LLM
Forgery Prompt Learner: To effectively convert forgery-related features into inputs for the LLM, we propose a Forgery Prompt Learner to transform the forgery segmentation map, the forgery score, and the consistency maps into forgery prompt embeddings. At the same time, we add a learnable prompt embeddings for the forgery prompt learner to incorporate extra information for the deepfake detection task. The Forgery Prompt Learner consists of two convolutional neural networks, one fully connected layer, and the learnable prompt embeddings $E_{\emph{base}}\in\mathbb{R}^{n_{1}\times C_{\emph{emb}}}$, where $C_{\emph{emb}}$ represents the dimensionality of the embedding vectors. Specifically, the two convolutional networks transform the forgery segmentation map and consistency maps into vector representations, $E_{\emph{loc}}\in\mathbb{R}^{n_{2}\times C_{\emph{emb}}}$ and $E_{\emph{cons}}\in\mathbb{R}^{n_{3}\times C_{\emph{emb}}}$, respectively. The forgery score is expanded into $E_{\emph{cls}}\in\mathbb{R}^{1\times C_{\emph{emb}}}$. These embeddings are concatenated and fed into the convolution layer to generate the forgery prompt embeddings $E_{\emph{forgery}}\in\mathbb{R}^{n_{f}\times C_{\emph{emb}}}$. Finally, the forgery prompt embeddings, visual prompt embeddings, and question prompt embeddings are input into the LLM.
LLM: The LLM processes the prompt embeddings to interpret the context and accurately identify forged regions. By integrating visual details (from $E_{\emph{forgery}}$ and $E_{\emph{visual}}$) with user queries (from $E_{\emph{question}}$), the LLM produces responses that provide forgery detection judgments and precisely localize manipulated regions (e.g., eyes, mouth). Here, we employ prompt tuning and LoRA to fine-tune the LLM using simulated image-text pairs specifically tailored for deepfake detection tasks. To ensure the LLM generates accurate responses, we use cross-entropy loss to quantify the discrepancy between the predicted response and the target label. The formula is defined as follows:
@@
where $\hat{y_{j}}$ represents the predicted probability for the $j$-th token, and $y_{j}$ denotes the corresponding ground truth label.
@@ 3.3 Data for LLM Prompt Tuning
Forgery Data Simulation: We intend for the LLM to identify pristine and deepfake images, while also locating the forged regions. This requires training on image-text pairs specifically depicting manipulated areas, which are currently unavailable. To address this gap, we draw on techniques from SBI (Shiohara & Yamasaki, 2022) to construct image-text pairs using existing real images. First, we generate facial landmarks from a real image $I_{\emph{real}}$, then randomly select 1 to $n$ regions (e.g., the nose, mouth, or eyes) as target forgery areas. We apply a slight affine transformation to the real image, resulting in an affine-transformed image $I_{\emph{affine}}$. The original real image is used as the background (target face), and the affine-transformed image serves as the foreground (source face). Following the approach in (Nguyen et al., 2024), we apply Poisson blending to combine the foreground and background images. The blending process is defined as follows:
@@
Methods Intra-dataset Cross-dataset FF++ CDF2 DFD DFDC CDF1 DFDCP AUC AP AUC AP AUC AP AUC AP AUC AP AUC AP Xception (Rossler et al., 2019) CVPR’19 97.23 96.44 81.65 88.74 89.75 - - - 80.98 90.00 69.90 81.95 FaceXRay+BI (Li et al., 2020a) CVPR’20 - - 79.50 - 95.40 93.34 - - 80.58 73.33 80.92 72.65 F3Net (Qian et al., 2020) ECCV’20 98.20 96.17 78.88 86.23 93.33 97.70 71.77 71.99 81.11 89.61 73.50 79.30 Multiattentional (Zhao et al., 2021a) CVPR’21 - - 68.26 - 92.95 - - - - - 63.02 - SPSL (Liu et al., 2021) CVPR’21 96.91 89.37 79.86 87.83 96.12 98.20 66.16 71.13 85.02 92.17 75.86 82.64 PCL+I2G (Zhao et al., 2021b) CVPR’21 99.11 - 90.03 - 99.07 - - - - - 74.27 - RECCE (Cao et al., 2022) CVPR’22 99.32 97.25 82.31 88.26 98.26 98.40 69.58 69.98 81.49 89.04 71.49 77.04 SBI (Shiohara & Yamasaki, 2022) CVPR’22 99.15 99.15 93.82 92.99 96.32 96.13 74.47 74.09 93.44 93.77 90.95 85.98 SFDG (Wang et al., 2023a) CVPR’23 - - 75.83 - 88.00 - - - - - 73.63 - AltFreezing (Wang et al., 2023b) CVPR’23 93.81 98.74 89.50 87.41 98.50 94.59 64.75 67.52 88.48 92.79 64.05 76.22 CADDM (Dong et al., 2023) CVPR’23 99.70 - 93.88 - 99.03 - 73.85 - 85.68 - 74.19 - UCF (Yan et al., 2023) CVPR’23 98.69 97.99 83.73 90.10 94.50 98.04 75.11 74.76 86.08 91.78 80.50 77.16 TALL (Xu et al., 2023) ICCV’23 99.87 - 90.79 - - - 76.78 - - - - - LSDA (Yan et al., 2024) CVPR’24 - - 91.10 - - - 77.00 - - - - - LAANet (Nguyen et al., 2024) CVPR’24 99.96 - 95.40 97.64 99.51 99.40 - - - - 86.94 97.70 CFM (Luo et al., 2024) TIFS’24 - - 89.65 - - - 80.22 - - - - - ED (Ba et al., 2024) AAAI’24 - - 93.6 - - - 75.4 - - - 90.2 - ProDet (Cheng et al., 2024) NIPS’25 - - 92.62 96.05 - - 71.52 72.8 94.48 96.66 82.83 88.89 RepDFD (Lin et al., 2025) NIPS’25 - - 89.94 - - - 80.99 - - - 95.03 - UDD (Fu et al., 2025) AAAI’25 - - 93.13 - - - 81.21 - - - 88.11 - Ours 99.53 99.54 94.71 93.59 99.64 99.68 79.12 77.69 97.62 97.67 91.81 88.26
@@
Table 1: Intra-dataset and cross-dataset evaluation of KFD compared with existing deepfake detection methods. The highest AUC and AP scores among all methods are highlighted in bold, while the second-highest scores are underlined. All experimental results are sourced from the original papers or reproduced using publicly available code repositories.
@@
Question and Answer Content: Training an LVLM requires a large number of visual question-answer pairs. Therefore, we construct corresponding text queries for each image. To ensure compatibility with the deepfake detection task, we first include a background description in each query, for example: “This is a facial image designed for deepfake detection, and it should not exhibit any localized color discrepancies or evident signs of splicing.”, which can be regarded as a kind of human prior knowledge. Additionally, we incorporate the prediction results from the Knowledge-guided Forgery Detector (KFD) into the prompt, such as “According to KFD prediction, the forgery score is 0.93.” We then ask a question related to the content of the image, such as “Is this a deepfake image?” The LVLM’s response states whether a forgery is present in the image and where the forgery area is. For instance, “Yes. This is a deepfake image, and the artifact is at the center face of the image.” Here, the forgery regions are defined according to those selected during forgery data simulation. By defining both queries and responses, we can train the LVLM to distinguish between pristine and deepfake images. The prompt format input to the LVLM follows this format:
@@
where $E_{visual}$ represents visual prompt embeddings, $E_{{forgery}}$ denotes the forgery prompt embeddings learned by the Forgery Prompt Learner, KFD result indicates the forgery classification results, and Task description provides a textual description of deepfake detection task.